OpenStudy (anonymous):

Find the altitude of a zone whose area is 240pi sq. u. if the circumference of a great circle is 24pi units.

OpenStudy (elonasushchik):

I think you mean "latitude" instead of "altitude." I think you are asking about the area of a region bounded by two lines of latitude. This appears to be a mathematics question. If the circumference of a great circle is 2πR = 24π units then the radius of the circle is R = 12 units. The surface area of the sphere is 4πR^2 = 4*π*12^2 = 576π sq. units. At a latitude λ the circumference of a line of latitude is Ccosλ where C = 24π is the circumference of a great circle. The area of a narrow strip of width dS at latitude λ is dA = CcosλdS. But dS = Rdλ, so dA = Ccosλ.Rdλ = 24π.12.cosλdλ = 288π cosλdλ. The area between lines of latitude λ1 and λ2 is the integral of this expression, viz. A = 288π(sinλ2-sinλ1). This area must equal 240π, so sinλ2-sinλ1 = 240/288 = 5/6. To solve the problem we need to one one of the two boundary lines of latitude. Possibly you are assuming that they are symmetrical about the equator, ie λ1 = -λ2. If so then 2sinλ2 = 5/6 sinλ2 = 5/12 = 0.416667 sinλ2 = 0.429775 radians = 24.6°. ie the area between latitude lines -24.6° and +24.6° is 240π sq. units.

OpenStudy (elonasushchik):

Don't ask me what those boxes are all about?!

OpenStudy (anonymous):

Its alitude , i need to find the height